Circular No. 8464

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

     URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005D IN UGC 3856

Further to IAUC 8459, J. Burket and W. Li report the LOSS discovery, on an unfiltered KAIT image taken on Jan. 13.35 UT, of an apparent supernova (mag ∼ 18.4) located at α = 7^{h}26^{m}57^{s}.36, δ = +20^{o}22′53′′.4 (equinox 2000.0), which is 4′′.8 east and 37′′.8 north of the nucleus of UGC 3856. Additional KAIT magnitudes: 2004 Dec. 12.39, [20.5; 17.37, 19.4; 19.38, 19.9; 20.33, 19.7; 24.35, 19.4.

SUPERNOVAE 2004ha–2004hs

S. Blondin, European Southern Observatory; and J. L. Prieto, Ohio State University, on behalf of the ESSENCE project (cf. IAUC 8427), report the discovery of 19 supernovae on multiple R- and I-band images taken at the Blanco 4-m telescope by project members. Spectroscopy was obtained with the Gemini North and South (+ GMOS), Keck I (+ LRIS), Magellan Baade (+ IMACS), and Magellan Clay (+ LDSS2) telescopes, showing each to be of type Ia except 2004ho (which is of type II).

SN  2004 UT  α_{2000}  δ_{2000}  R  z
2004ha Nov. 6 2^{h}04^{m}27.^{s}00 − 4^{o}52′46′′.0 23.0 0.5
2004hb Nov. 6 2 08 09.70 − 4 41 51.8 23.4 0.7
2004hc Nov. 12 23 24 32.67 − 8 41 03.5 23.3 0.7
2004hd Nov. 14 2 08 48.21 − 4 26 10.4 22.5 0.5
2004he Nov. 14 2 29 48.79 − 8 20 45.9 23.0 0.6
2004hf Nov. 16 2 32 00.14 − 8 42 23.8 22.0 0.4
2004hg Nov. 16 2 34 55.19 − 8 30 43.6 22.7 0.6
2004hh Nov. 20 2 06 25.01 − 4 38 04.0 21.4 0.2
2004hi Nov. 20 2 08 38.82 − 5 08 11.7 22.4 0.4
2004hj Nov. 20 2 29 41.93 − 8 43 49.4 21.5 0.4
2004hk Dec. 5 23 27 04.38 − 8 38 45.1 21.5 0.3
2004hl Dec. 7 1 13 38.17 − 0 27 39.0 22.0 0.3
2004hm Dec. 7 2 28 03.11 − 7 42 29.7 21.5 0.2
2004hn Dec. 9 1 13 32.38 + 0 37 15.3 23.5 0.6
2004ho Dec. 9 2 06 16.03 − 3 52 27.7 22.3 0.3
2004hp Dec. 9 2 09 35.52 − 3 46 23.5 22.0 0.5
2004hq Dec. 9 2 30 18.04 − 8 22 25.0 23.1 0.7
2004hr Dec. 11 1 08 48.33 + 0 00 49.5 22.6 0.4
2004hs Dec. 13 2 09 33.69 − 4 13 04.0 23.1 0.4

2005 January 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green